OpenStudy (anonymous):

Find dy/dx of e^(2x)=sin(x+4y)

OpenStudy (solomonzelman):

well, the derivative of both sides will need the chain rule, and the derivative of 4y (which will be in the chain rule) will have a chain rule of (multiplying by) dy/dx.

OpenStudy (solomonzelman):

then as you differentiate, isolate the dy/dx by itself to find the answer.

OpenStudy (solomonzelman):

lets take it piece by piece. What is the derivative of \(\large\color{black}{ \displaystyle e^{2x} }\) ?

OpenStudy (anonymous):

2e^(2x)

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

What is the derivative of sin(x) ?

OpenStudy (anonymous):

cos(x)

OpenStudy (solomonzelman):

yes, but when it is sin(x+4y), (when the inner argument is x+4y) we will need the chain rule.

OpenStudy (solomonzelman):

what is the derivative of x+4y (derivatiev dy/dx) ?

OpenStudy (anonymous):

1+4dy/dx? I think this is where I am messing up

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

So, sin(x+4y) comes as follows, \(\large\color{black}{ \displaystyle \cos(x+4y)\times\left[1+4\frac{dy}{dx}\right] }\)

OpenStudy (solomonzelman):

saying we differnetiate the entire thing as if it is just sin(x), and then applied the chain rule

OpenStudy (anonymous):

cos(x+4y)*(1+4dy/dx)! I didn't use parenthesis for the )1+4dy/dx) I think that was my error

OpenStudy (solomonzelman):

So, from both sides, \(\large\color{black}{ \displaystyle e^{2x}=\sin(x+4y) }\) differentiating, \(\large\color{black}{ \displaystyle 2e^{2x}=\cos(x+4y)\times\left[1+4\frac{dy}{dx}\right] }\)

OpenStudy (solomonzelman):

yes, parenthesis are very important

OpenStudy (solomonzelman):

See how we differentiated the function ?

OpenStudy (anonymous):

Yep! Can you work me through setting the side equal to dy/dx just for practice?

OpenStudy (solomonzelman):

sure, you will need to expand the right side at first.

OpenStudy (solomonzelman):

can you do that for me?

OpenStudy (anonymous):

cos(x)+cos(4y)*(1+4dy/dx)? Or no? Rusty at math sorry I know this should be the simple part!

OpenStudy (solomonzelman):

where does cos(x) come from?

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 2e^{2x}=\cos(x+4y)\times\left[1+4\frac{dy}{dx}\right] }\) into, \(\large\color{black}{ \displaystyle 2e^{2x}=\cos(x+4y)\times\left[1\right] +\cos(x+4y)\times\left[4\frac{dy}{dx}\right]}\)

OpenStudy (solomonzelman):

\[A(b+C)=Ab+AC\]

OpenStudy (solomonzelman):

(same here)

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 2e^{2x}=\cos(x+4y)\times\left[1\right] +\cos(x+4y)\times\left[4\frac{dy}{dx}\right]}\) \(\large\color{black}{ \displaystyle 2e^{2x}=\cos(x+4y)+4\cos(x+4y)\times\left[\frac{dy}{dx}\right]}\)

OpenStudy (solomonzelman):

then solve for dy/dx.... can you go on?

OpenStudy (anonymous):

Just subtract them over right? Hahah

OpenStudy (solomonzelman):

yes, subtract the cos(x+4y) from both sides, and then divide both sides by 4cos(x+4y)

OpenStudy (anonymous):

Thank you so much!

OpenStudy (solomonzelman):

sure, yw!